OpenStudy (legomyego180):

The base of a solid is the circle x^2 + y^2 =16. Find the volume of the solid given that the cross sections perpendicular to the y − axis are squares.

OpenStudy (legomyego180):

How do I find my bounds / limits here?

OpenStudy (legomyego180):

@fwizbang @agent0smith

OpenStudy (legomyego180):

i know r=4 and the center is at the orgin

OpenStudy (fwizbang):

Is this meant to be a single double or triple integration?(i.e. what class are you in?)

OpenStudy (legomyego180):

calc 2 single int

OpenStudy (fwizbang):

OK. Since we know the cross section perpendicular to the y-axis, y is the variable to integrate. So, for a given y, what's the biggest and smallest x can get?

OpenStudy (legomyego180):

4

OpenStudy (fwizbang):

If y is 3, then x can't be 4 if x^2+y^2 < 16.....

OpenStudy (fwizbang):

Assume y is fixed and solve for x.

OpenStudy (legomyego180):

\[ysqrt{16-x^2}\]

OpenStudy (legomyego180):

Sorry having some technical difficulties

OpenStudy (fwizbang):

Good(I think). Remember that there's a positive and negative root so that\[-\sqrt{16-y^2}<x<sqrt{16-y^2\]. So this gives us the length of one side of the square.

OpenStudy (legomyego180):

\[y < \sqrt{16-x^2}\]

OpenStudy (fwizbang):

We're going to integrate y, so we want to solve for x.

OpenStudy (legomyego180):

Its just the same though isnt it?

OpenStudy (fwizbang):

In terms of the circle, yes. But the square cross sections are in the x-z plane,perpendicular to the y-axis, so it'll be easier to integrate over y.

OpenStudy (legomyego180):

So visually this is a tipped over cylinder with a bunch of squares stacked inside in three dimensions

OpenStudy (legomyego180):

How would you solve for x?

OpenStudy (legomyego180):

Sorry, I had a hard time grasping this concept in class too.

OpenStudy (fwizbang):

Since the edge of the circle is x^2+y^2=16, the interior is \[-\sqrt{16-y^2}<x< \sqrt{16-y^2}\]

OpenStudy (fwizbang):

I guess I visualized this with the y-axis pointing up and the x-z plane horizontal. Then the Hemisphere is essentially a stack of square plates.

OpenStudy (fwizbang):

So assuming y is fixed, what is the length of the square edge?

OpenStudy (legomyego180):

im not sure

OpenStudy (legomyego180):